# An Exploration of Default Images in Text-to-Image Generation

### v5

@@ 6.3. Explainability and Alignment of Text-to-Image Generation Models
daras2022 experimented with how gibberish prompts are understood by the model and suspected that TTI models have developed a ‘secret language’ (daras2022). Their hypothesis was that DALL-E 2 had made ‘hidden’ associations between certain terms in text prompts and concepts in images, such as ‘Apoploe vesrreaitais’ meaning ‘bird’. daras2022 found that these terms often produced consistent results in isolation but also sometimes in combination with other terms in prompts. Our work is similar in that some input prompts reliably sample from the same region in latent space. We found that with prompts that include both known and unknown terms, the model often generates images from known terms, omitting unknown terms. However, when Midjourney does not ignore unknown terms, it may generate images sharing visual motifs with default images. However, unlike in a ‘secret language’, this is not a one-to-one mapping between prompt and resulting image. In our case, several unrelated prompts can produce variations of the same default image, as demonstrated in Figure 2.
Because it is often unclear why the model produced a certain default image, the problem of default images touches on research in interpretability and Explainable AI (XAI). Default images reflect inherent biases and patterns encoded in the TTI model. In a sense, the default image is a nonce output and this output currently is the only mechanism of the TTI system to let the user know that the prompt is not understood. Other generative models, such as language models, include clear, explicit ways of letting users know that something is not understood or that they cannot respond (e.g., by refusing to answer for reasons of content policy). TTI models lack these mechanisms by default. In practice, language models internally handle such content and safety checks on Midjourney. Our prompts, however, pass these checks and obscure the fact that the model can not respond to the user prompt as expected. Instead of returning nothing, the TTI model generates default images. This could be termed a case of ostensive inferential communication (sperber1995relevance; wilson2004relevance), a model of communication where a speaker provides cues (ostension) to make their intended meaning clear, and the listener infers that meaning through a process of reasoning (inference). This miscommunication is also a plausible explanation as to why user satisfaction is lower when the image is different from expectations.
@@ 6.4. The Future of Default Images
However, the phenomenon of default images itself will likely not disappear entirely. We will continue to see situations where words in prompts are not understood by the generative model, where users type incoherent prompts (11084610) in error, or where users intentionally experiment with unfamiliar language, unusual phrasing, or invented terms to push the model beyond known outputs. In such cases, the model may revert to visual patterns that function as a “fallback,” even if these patterns become less recognizable as defaults in future generative systems.
@@ 6.6. Limitations and Future Work
Our focus on Midjourney limits the external validity of findings to other TTI systems. We selected this scope to allow in-depth qualitative and quantitative analysis within a consistent model environment. Future research could adopt a comparative lens, investigating default images in other TTI generation systems, such as Stable Diffusion (Rombach_2022_CVPR), Imagen (2205.11487.pdf), Flux (labs2025flux1kontextflowmatching), Janus (janus), or video generation models, potentially revealing differences in model capabilities and biases beyond canonical benchmarks.
Due to the vastness of Midjourney’s latent space, our analysis cannot be exhaustive (2507.17922.pdf; 3491102.3501825.pdf). Each model version introduces its own set of default images, with some occurring far more frequently than others (cf. Figure 6). As a result, our work can only provide a partial view of the phenomenon. Nonetheless, certain defaults (like the Lady-Birdhead image), are common and straightforward to identify even in small-scale manual studies, suggesting that foundational insights can be gained without full coverage. Future work could extend our approach to systematically map default images across different model versions. Default images could form the basis of a benchmark for model evaluation, capturing how systems respond to unknown or ambiguous inputs. Such a benchmark would complement existing technical metrics by foregrounding model behavior in edge-case scenarios encountered in creative practice.
@@
We did not directly investigate bias in default images in this work, yet their occurrence in the absence of explicit text guidance potentially provides an opportunity to study “raw” biases in TTI models. Building on related studies on bias in TTI models (e.g., (Vasilev2024S137; Elsharif2025122636; Saravanan202384; Kandwal202473; luccioni2023stablebiasanalyzingsocietal; 10.1145/3706599.3719678; jha2024visageglobalscaleanalysisvisual; vice2023quantifyingbiastexttoimagegenerative)), future research could analyze demographic, cultural, or aesthetic biases embedded in default images.
Studying default images from a color theory perspective could also provide new insights on prevailing harmonies and contrasts and contribute to an understanding of the symbolism and potential impact of such associations. Exploring how default images tie into the creative practice could be an interesting avenue for future studies. Future work could examine the implications of default images for creativity support, for example through Construal Level Theory (Trope2010-TROCTO) or other psychological frameworks.
@@ Appendix B Dataset
Reactions Reactions to the message from Discord users $\bigstar$ (1)

### v6

@@ 6.3. Explainability and Alignment of Text-to-Image Generation Models
Daras and Dimakis experimented with how gibberish prompts are understood by the model and suspected that TTI models have developed a ‘secret language’ (Daras and Dimakis, 2022). Their hypothesis was that DALL-E 2 had made ‘hidden’ associations between certain terms in text prompts and concepts in images, such as ‘Apoploe vesrreaitais’ meaning ‘bird’. Daras and Dimakis found that these terms often produced consistent results in isolation but also sometimes in combination with other terms in prompts. Our work is similar in that some input prompts reliably sample from the same region in latent space. We found that with prompts that include both known and unknown terms, the model often generates images from known terms, omitting unknown terms. However, when Midjourney does not ignore unknown terms, it may generate images sharing visual motifs with default images. However, unlike in a ‘secret language’, this is not a one-to-one mapping between prompt and resulting image. In our case, several unrelated prompts can produce variations of the same default image, as demonstrated in Figure 2.
Because it is often unclear why the model produced a certain default image, the problem of default images touches on research in interpretability and Explainable AI (XAI). Default images reflect inherent biases and patterns encoded in the TTI model. In a sense, the default image is a nonce output and this output currently is the only mechanism of the TTI system to let the user know that the prompt is not understood. Other generative models, such as language models, include clear, explicit ways of letting users know that something is not understood or that they cannot respond (e.g., by refusing to answer for reasons of content policy). TTI models lack these mechanisms by default. In practice, language models internally handle such content and safety checks on Midjourney. Our prompts, however, pass these checks and obscure the fact that the model can not respond to the user prompt as expected. Instead of returning nothing, the TTI model generates default images. This could be termed a case of ostensive inferential communication (Sperber and Wilson, 1995; Wilson and Sperber, 2006), a model of communication where a speaker provides cues (ostension) to make their intended meaning clear, and the listener infers that meaning through a process of reasoning (inference). This miscommunication is also a plausible explanation as to why user satisfaction is lower when the image is different from expectations.
@@ 6.4. The Future of Default Images
However, the phenomenon of default images itself will likely not disappear entirely. We will continue to see situations where words in prompts are not understood by the generative model, where users type incoherent prompts (Kuhn et al., 2025) in error, or where users intentionally experiment with unfamiliar language, unusual phrasing, or invented terms to push the model beyond known outputs. In such cases, the model may revert to visual patterns that function as a “fallback,” even if these patterns become less recognizable as defaults in future generative systems.
@@ 6.6. Limitations and Future Work
Our focus on Midjourney limits the external validity of findings to other TTI systems. We selected this scope to allow in-depth qualitative and quantitative analysis within a consistent model environment. Future research could adopt a comparative lens, investigating default images in other TTI generation systems, such as Stable Diffusion (Rombach et al., 2022), Imagen (Saharia et al., 2022), Flux (Labs et al., 2025), Janus (Chen et al., 2025), or video generation models, potentially revealing differences in model capabilities and biases beyond canonical benchmarks.
Due to the vastness of Midjourney’s latent space, our analysis cannot be exhaustive (Quaye et al., 2025; Liu and Chilton, 2022b). Each model version introduces its own set of default images, with some occurring far more frequently than others (cf. Figure 6). As a result, our work can only provide a partial view of the phenomenon. Nonetheless, certain defaults (like the Lady-Birdhead image), are common and straightforward to identify even in small-scale manual studies, suggesting that foundational insights can be gained without full coverage. Future work could extend our approach to systematically map default images across different model versions. Default images could form the basis of a benchmark for model evaluation, capturing how systems respond to unknown or ambiguous inputs. Such a benchmark would complement existing technical metrics by foregrounding model behavior in edge-case scenarios encountered in creative practice.
@@
We did not directly investigate bias in default images in this work, yet their occurrence in the absence of explicit text guidance potentially provides an opportunity to study “raw” biases in TTI models. Building on related studies on bias in TTI models (e.g., (Vasilev et al., 2024; Elsharif et al., 2025; Saravanan et al., 2023; Kandwal and Nehra, 2024; Luccioni et al., 2023; Lan et al., 2025; Jha et al., 2024; Vice et al., 2025)), future research could analyze demographic, cultural, or aesthetic biases embedded in default images.
Studying default images from a color theory perspective could also provide new insights on prevailing harmonies and contrasts and contribute to an understanding of the symbolism and potential impact of such associations. Exploring how default images tie into the creative practice could be an interesting avenue for future studies. Future work could examine the implications of default images for creativity support, for example through Construal Level Theory (Trope and Liberman, 2010) or other psychological frameworks.
@@ 7. Conclusion
Acknowledgements.
This work is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) Vanier Scholarship CGV-186947. The last author gratefully acknowledges support from the University of Oulu’s Centre for Applied Computing (CAC) and the NOSTE development programme for education.
@@ Appendix B Dataset
Reactions Reactions to the message from Discord users $\ast$ (1)